he demand and supply for a particular commodity are given by the following two equations:

Demand: P = 10 – 0.2Qd

and

Supply: P = 2 + 0.2Qs

Where Qd and Qs are quantity demanded and quantity supplied, respectively, and P is price.

Using the equilibrium condition Qs = Qd, determine equilibrium price and equilibrium quantity.

Equilibrium price = $

Equilibrium quantity =  units

Graph the two equations to substantiate your answer.


Instructions:
1. Use the line tools Qd and Qs to draw the demand and supply curves for P = 4 and 8.
2. Use the drop line tool E to identify the equilibrium quantity and price.
Answer #1

P = 10-.2*Qd

Qd = (1/.2)*(10 - P)

Qd = 5*(10-P) -------------- (1)

P = 2 + 0.2*Qs

Qs = (5)*(P-2) ----------------(2)

At equilibrium.

Qd = Qs

5*(10-P) = (5)*(P-2)

10-P = P-2

2P = 12

P = 12/2

P = $6

Q = 5*(10-6)

Q = 20 units

So,

Equilibrium price = $6

Equilibrium quantity = 20 units

Price ($) Quantity demanded (5*(10-P)) Quantity supplied (5*(P-2))
4 30 10
8 10 30